OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -36 and 2304, respectively.

OpenStudy (welshfella):

the nth term of a G S is given by a1 r^(n-1) where a1 = first term and r = common difference

OpenStudy (welshfella):

so the second term = a1 r and 5th term = a1 r^4 so we have a1 r^4 2301 ---- = ---- a1 r -36

OpenStudy (welshfella):

Can you find the value of r?

OpenStudy (mathmale):

Basically correct. Might be clearer to draw the formula for arith. sequence, or to use Equation to make it fancy-clear.\[a _{n=}a _{0}*r ^{(n-1)}\]

OpenStudy (mathmale):

a sub 0 is the first value of the sequence. r is the common factor.

OpenStudy (mathmale):

When you come back online, please let the rest of us know what further info you may need to get started.

OpenStudy (anonymous):

i got it guys

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

Find the sum of the arithmetic sequence. -4, -1, 2, 5, 8, 11, 14 how do you do that?

OpenStudy (mathmale):

There's a simple formula for that. If s is the sum of the sequence, a is the first term, n is the number of terms, and d is the difference between terms, s = a + (n - 1)*d

OpenStudy (mathmale):

You've typed in a sequence. What are the values of a, n and d here?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

-4, 7, and i dont know the last one

OpenStudy (mathmale):

How far apart are -4 and -1? How far apart are 5 and 8? This should tell you the value of "d."

OpenStudy (anonymous):

three

OpenStudy (mathmale):

Right. So, a=-4, n=7 and d=3. Can you now find the sum of these 7 terms?

OpenStudy (mathmale):

You could check your result by actually adding them up.

OpenStudy (anonymous):

waitso i just plug it in

OpenStudy (anonymous):

i got 14 ??????

OpenStudy (anonymous):

nvmgot it, 35 right

OpenStudy (welshfella):

35 is correct